# Weiwei Ding1,2, Chuanguang Yang1,2, Yongjun Xu1,2, Xiaolong Jin1,2,3, Jiafeng Guo1,2,3 1

### v1

@@ A.4 Implementation Details
We use the Qwen2.5-Coder-7B-Base as our backbone LLM. For the cold-start stage, our implementation is based on the LlamaFactory framework (Zheng et al., 2024); we use a cosine learning rate schedule with an initial learning rate of 2e-5, a batch size of 8, and a maximum sequence length of 4096. We fine-tune the model for 8 epochs. In the RL stage, our implementation is based on Agent-R1 framework (Ouyang et al., 2025) (a verl-based (Sheng et al., 2024) agentic training framework); we set the number of rollouts to N=8, use a batch size of 64 for rollout, and a global batch size 512 for training. We train the model for 1 epoch. At validation, we set the temperature to 0 to ensure deterministic outputs. The detail parameters can be found in 5. All experiments were conducted on 8 $\times$ NVIDIA H100 (80GB) GPUs and 8 $\times$ NVIDIA A100 (80GB) GPUs.

### v2

@@ A.4 Implementation Details
We use the Qwen2.5-Coder-7B-Base as our backbone LLM. For the cold-start stage, our implementation is based on the LlamaFactory framework (zheng2024llamafactory); we use a cosine learning rate schedule with an initial learning rate of 2e-5, a batch size of 8, and a maximum sequence length of 4096. We fine-tune the model for 8 epochs. In the RL stage, our implementation is based on Agent-R1 framework (Agent-R1) (a verl-based (sheng2024hybridflow) agentic training framework); we set the number of rollouts to N=8, use a batch size of 64 for rollout, and a global batch size 512 for training. We train the model for 1 epoch. At validation, we set the temperature to 0 to ensure deterministic outputs. The detail parameters can be found in 5. All experiments were conducted on 8 $\times$ NVIDIA H100 (80GB) GPUs and 8 $\times$ NVIDIA A100 (80GB) GPUs.